[Speculative attack]

Assume Thailand unilaterally defends an exchange rate of 1 dollar to 24 baht with a foreign reserve of 40 billion US dollar. Suppose an investor wants to implement a speculative attack on baht.
(a)
What is the minimum amount of baht that this investor should sell so that the Thai government will
use all her foreign reserves to defend the exchange rate?

(b) Suppose the investor borrows just enough baht for the attack and the transaction cost for this borrowing is 800 million dollars. The investor sells all his baht in the foreign exchange market to buy US dollars (at the rate of 1 dollar to 24 baht). Suppose with probability 0.2 the attack is successful and the Thai government runs out of the foreign reserve. If the attack is successful, then the Thai government devalues baht and the new exchange rate would be 1 dollar to 40 baht. Given this information, what is the
investor’s expected profit from the speculative attack (in terms of US dollar)?
Answer #1

a)thai govt will use all her foreign reserves in order to defend exhange rate, which means 40 billion dollar will be used....

investor has to sell baht atleast equal to value of 40 billion dollar at rate of 1$ to 24 baht

40 billion dollar = 24* 40 billion baht = 960 billion baht

the minimum amount of baht that this investor should sell= 960 billion baht

b) investors expected profit =probability of successful* profit if successful + probability of failure *profit if not successful

profit if successful ; now in new excahange rate 1$ to 40 baht, by seeling 960 billion baht ,

investor will get= 960 billion baht/40 baht = 24 billion dollar

but by playing this attack , he got 40 billion dollar

so profit of 40-24 billion dollar = 16 billion dollar

after considering the borrwing cost , net profit = 16- 0.8 billion dollars = 15.2 billion dollar

if not succesful, then exchange rate will not change... net loss= borrowing cost= 0.8 billion dollar   

investors expected profit =probability of successful* profit if successful + probability of failure *profit if not successful

= 0.2*15.2 billion dollar + 0.8* (-0.8 billion dollar)

= 3.04 billion dollar - 0.64 billion dollar = 2.4 billion dollar